# A Visual Case Study of the Training Dynamics in Neural Networks

### v1

@@ Abstract
This paper introduces a visual sandbox designed to explore the training dynamics of a small-scale transformer model, with the embedding dimension constrained to $d=2$. This restriction allows for a comprehensive two-dimensional visualization of each layer’s dynamics. Through this approach, we gain insights into training dynamics, circuit transferability, and the causes of loss spikes, including those induced by the high curvature of normalization layers. We propose strategies to mitigate these spikes, demonstrating how good visualization facilitates the design of innovative ideas of practical interest. Additionally, we believe our sandbox could assist theoreticians in assessing essential training dynamics mechanisms and integrating them into future theories. The code is available at https://github.com/facebookresearch/pal.
@@ 1 Introduction
We believe that improving the understanding of this matter is of great importance not only to enhance performance but also to limit the failure cases that naturally occur when scaling the size of the models (Chowdhery et al., 2024; Dehghani et al., 2023; Molybog et al., 2023). This would notably contribute to the reduction of memory and budget costs, along with carbon emissions associated with training large models (Faiz et al., 2024). In this paper, we deliberately choose to focus on a small model to gain a deeper understanding of its training dynamics. In particular, the crux of our approach is to consider a model small enough to visualize it fully. Specifically, we focus on a transformer architecture with an embedding dimension $d=2$, which allows us to plot the dynamics of each layer in a two-dimensional plane. This approach enables us to analyze the model’s internal workings and visualize the underlying mechanisms that drive its behavior, which may be obscured in larger models. We hope that insights gained from small-scale studies could provide a foundation for understanding the complex dynamics at play in larger models, ultimately shedding light on their behavior and performance.
Our main contribution is to provide a visual sandbox which can be of great interest for several reasons. Theoreticians can use it to build intuition on key mechanisms underlying training dynamics, potentially leading to new theorems. Practitioners can use it to test new ideas, such as optimizer modifications. In particular, modifications to the architecture, to the training setting, or to the data can be seamlessly integrated into our pipeline. By providing this sandbox, we aim to facilitate the understanding of the complex dynamics at play in larger models.
Summary of our contributions.
Our contributions are as follows:
  • We develop a visual sandbox to visualize the training dynamics of transformers. Our code is made available to the reviewers and will be open-sourced. It is designed to create videos that help comprehend the model’s internal dynamics comprehensively.
  • Using our visual sandbox, we provide insights into the training dynamics of our model. We classify the types of circuits learned for our problem and illustrate the typical two-phase learning process, consisting of representation learning followed by classifier fitting.
  • We detail the transferability of circuits to showcase the usefulness of curriculum learning and data curation.
  • We investigate loss spikes, suggesting potential strategies for mitigation, which could lead to more stable training processes.
@@
Opening the Black Box. As neural networks increasingly permeate civil society, concerns about their opacity intensify. This has fueled the ongoing quest to "open the black box." To tackle this issue, researchers have developed various methods, such as extracting meaningful features from neural network activations (Fel et al., 2023), and assessing the impact of perturbations on model inputs (Koh & Liang, 2017; Fel et al., 2021), among others. Recently, the field of mechanistic interpretability has advocated for exposing the internal mechanisms of transformers to provide novel insights into their capabilities (Elhage et al., 2021; Olsson et al., 2022). While some methodologies may apply to large language models (Templeton et al., 2024), precise ablation studies often focus on controlled environments and simplified architectures (Rauker et al., 2023; Charton, 2022; Liu et al., 2022; Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022; Wang et al., 2023; Cabannes et al., 2024a). Our work aligns with this field, as we aim to make the internal behavior of transformers more explicit through carefully selected visualizations in controlled settings. In particular, by utilizing low-dimensional embeddings, we effectively circumvent the curse of dimensionality as defined by (Olah, 2023). Although low-dimensional embeddings make learning the task more challenging, they enable comprehensive visualization of every model component. However, unlike mechanistic interpretability, our goal is to provide insights into the training dynamics of these models, in the spirit of Wortsman et al. (2024).
@@
Training Dynamics of Neural Networks. With the advent of deep learning, new optimization challenges have emerged. Training loss functions are no longer necessarily convex or smooth, the optimization landscape has expanded enormously, and computations are distributed to manage this scale. As the scale of models increases, the absence of theoretical results and closed-form solutions for optimizing neural networks underscores the need to better understand their behavior in practice during training, which is increasingly becoming a computational bottleneck. One approach to address this issue is to employ mathematical abstractions, as seen with neural tangent kernels (NTK) Jacot et al. (2018); Chizat et al. (2020) and mean-field analysis (Mei et al., 2018; Chen et al., 2024), among others (Abbe et al., 2022; Ahn et al., 2024; Cabannes et al., 2024b). Unfortunately, obtaining formal, rigorous results often requires simplifications that deviate from practical implementations. This type of analysis can overlook crucial details that significantly alter the training dynamics.
2 Sandbox Setup
@@
In this section, we describe our controlled setup, focusing on both the data and the architecture.
@@ 2.1 Sparse Modular Addition
To probe our model’s decision-making process, we concentrate on a straightforward mathematical task. We draw inspiration from the sparse parity problem (Barak et al., 2023; Abbe et al., 2023) and the modular addition problem (Power et al., 2022; Nanda et al., 2023) to formulate the sparse modular addition problem. This problem is characterized by the following parameters:
@@
$\begin{array}[]{cccc}f^{*}:&\mathbb{F}_{p}^{L}&\to&\mathbb{F}_{p}\\ &x=(x_{1},x_{2},\ldots,x_{L})&\mapsto&y=\sum_{t\in[k]}x_{t}.\end{array}$
@@
The sparse modular addition task offers the benefits of simplicity while allowing for the observation of a variety of training behaviors.
2.2 One-layer Transformer Architecture
This paper focuses on a specific architecture designed to address our problem.
@@ 2.2 One-layer Transformer Architecture
The sparse modular addition problem is inherently discrete. To handle it with a differentiable architecture, we need to embed it in a continuous space ${\mathbb{R}}^{d}$. We first embed each token with both semantic and positional information. Given a learnable token embedding $E:{\mathbb{F}}_{p}\to{\mathbb{R}}^{d}$, and a learnable position embedding $P:[L]\to{\mathbb{R}}^{d}$, a sentence in token space is lifted to a sentence in embedding space through the following embedding operation
$\forall\,t\in[L],\qquad z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}.$ (embedding)
This type of embedding, known as absolute position embedding, incorporates both the semantic meaning of the token and its position within the sequence. The normalization, which projects each element of the sentence onto the sphere, is known as RMS-norm (Zhang & Sennrich, 2019).
Next, an attention mechanism is applied to the sequence of normalized embeddings $(z_{t})$ to aggregate them into a single sentence embedding $\xi\in{\mathbb{R}}^{d}$. It utilizes a query vector $q\in{\mathbb{R}}^{d}$, and a value matrix $V\in{\mathbb{R}}^{d\times d}$. Denoting $z=(z_{t})\in{\mathbb{R}}^{d\times L}$, the sentence embedding can be expressed in matrix form as
@@
where the softmax operation is defined as a mapping from ${\mathbb{R}}^{L}$ to the simplex $\Delta_{L}$ in ${\mathbb{R}}^{L}$. Specifically, the $s$-th component of $\operatorname{softmax}((a_{t}))$ is proportional to $\exp(a_{s})$, formally expressed as $\operatorname{softmax}((a_{t})_{t\in[L]})=(\exp(a_{s})/\sum_{t\in[L]}\exp(a_{t}))_{s\in[L]}$. Compared to the attention mechanism in Vaswani et al. (2023), we omit both the key and output matrices, which would act as extra parameters that do not increase the expressivity of our model.
@@
$\zeta:=\xi+\sum_{i\in[h]}\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}+b_{i}\right)\cdot u_{i}\in\mathbb{R}^{d},$
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a non-linear function, chosen to be $\sigma(x)=x\varphi(x)$ with $\varphi$ being the Gaussian cumulative function. This function $\sigma$ is known as the GELU activation. The final embedding vector $\zeta$ is decoded back to token space based on how it aligns with the respective token embeddings. A softmax layer converts these alignments into a probability vector over the different output classes:
$p_{\zeta}(y)=\operatorname{softmax}((E(y)^{\top}\zeta)_{y\in\mathbb{F}_{p}}).$
Abstracting all the learnable weights into a single parameter $\theta$, our architecture takes as input a sentence $x\in\mathbb{F}_{p}^{L}$ and outputs a probability vector over the classes $p_{\theta}(y|x)$. The parameters of our model are optimized by minimizing the cross-entropy loss defined as:
$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[-\log(p_{\theta}(Y|X))\right],$
where $p$ is a training distribution, typically chosen as the counting measure over some training data. This loss is a proxy for the measure we aim to optimize, which is accuracy, defined as:
$\mathcal{L}(\theta):=\mathbb{E}_{(X,Y)\sim p}\left[\mathbf{1}_{\{Y=\operatorname*{arg\,max}_{y}p_{\theta}(y|X)\}}\right],$
with $p$ denoting a data distribution, typically chosen as the counting measure over some testing data.
Default Configuration.
In our experiments, we trained our networks using $n=2048=2^{11}$ data points, which were sampled uniformly with replacement from the $p^{L}$ possible sentences. We utilized the Adam optimizer with parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$ (Kingma & Ba, 2017), and we initialized the network weights using the default schemes provided by PyTorch (Paszke et al., 2019).
2.3 Visualization Tools
The primary motivation of this paper is to embed all computations in dimension $d=2$, enabling us to visualize everything occurring within our model in a relatively straightforward manner. Our codebase is designed to generate videos of the training dynamics, tracking several key aspects. We notably track the following.
@@ 2.3 Visualization Tools
Transform level lines. We visualize the mapping from sentence embedding $\xi$ to their associated learned probabilities $p_{\zeta}(y=1)$. We also plot the sentence embedding on the same plot to better understand the level line changes.

### v2

# Clustering Head: A Visual Case Study of the Training Dynamics in Transformers

@@ Abstract
This paper introduces the sparse modular addition task and examines how transformers learn it. We focus on transformers with embeddings in ${\mathbb{R}}^{2}$ and introduce a visual sandbox that provides comprehensive visualizations of each layer throughout the training process. We reveal a type of circuit, called "clustering heads," which learns the problem’s invariants. We analyze the training dynamics of these circuits, highlighting two-stage learning, loss spikes due to high curvature or normalization layers, and the effects of initialization and curriculum learning.
Mechanistic Interpretability, Circuits, Clustering Heads, Transformers
[…]
@@ 1 Introduction
In this paper, we have deliberately chosen to focus on a small model to gain a deeper understanding of its training dynamics. In particular, the crux of our approach is to consider a model small enough to visualize it fully. Specifically, we first focus on a transformer architecture with an embedding dimension $d=2$, which allows us to plot the dynamics of each layer in a two-dimensional plane. We then generalize our insights to larger embedding dimensions. This approach enables us to analyze the model’s internal workings and visualize the underlying mechanisms that drive its behavior, which may be obscured in larger models. We hope that insights gained from small-scale studies could provide a foundation for understanding the complex dynamics at play in larger models, ultimately shedding light on their behavior and performance.
Our contributions can be summarized as follows:
  • We introduce the sparse modular addition problem and develop a visual sandbox to visualize the training dynamics of transformers. Our code is available at https://github.com/facebookresearch/pal. It is designed to create videos that help thoroughly comprehend the model’s internal dynamics.
  • Using our visual sandbox, we unveil Clustering Heads, circuits that capture the invariants of the task at hand, and we classify concrete realizations that can be seen as variations of these idealized circuits.
  • We study the training dynamics behind clustering heads, illustrating a two-phase learning process consisting of representation learning followed by classifier fitting. We investigate loss spikes, suggesting potential strategies for mitigation, which could lead to more stable training processes. We also peek into the transferability of circuits to showcase the usefulness of curriculum learning and data curation.
  • Our theoretical insights show that clustering heads can be learned via gradient descent and provide an understanding of the two-stage learning observed in practice.
@@
Opening the Black Box. As neural networks increasingly permeate civil society, concerns about their opacity intensify. This has fueled the ongoing quest to "open the black box." To tackle this issue, researchers have developed various methods, such as extracting meaningful features from neural network activations (Fel et al., 2023), and assessing the impact of perturbations on model inputs (Koh & Liang, 2017; Fel et al., 2021), among others. Recently, the field of mechanistic interpretability has advocated for exposing the internal mechanisms of transformers to provide novel insights into their capabilities (Elhage et al., 2021; Olsson et al., 2022). While some methodologies may apply to large language models (Templeton et al., 2024), precise ablation studies often focus on controlled environments and simplified architectures (Rauker et al., 2023; Charton, 2022; Liu et al., 2022; Nanda et al., 2023; Geva et al., 2023; Meng et al., 2022; Wang et al., 2023; Cabannes et al., 2024a). Our work aligns with this field, since we aim to make the internal behavior of transformers more explicit through carefully selected visualizations in controlled settings. In particular, by utilizing low-dimensional embeddings, we effectively circumvent the curse of dimensionality as defined by Olah (2023). Although low-dimensional embeddings make learning the task more challenging, they enable comprehensive visualization of every model component. However, unlike mechanistic interpretability, our goal is to provide insights into the training dynamics of these models, in the spirit of Wortsman et al. (2024).
@@
Training Dynamics of Neural Networks. With the advent of deep learning, new optimization challenges have emerged. Training loss functions are no longer necessarily convex or smooth, the optimization landscape has expanded enormously, and computations are distributed to manage this scale. As the scale of models increases, the absence of theoretical results and closed-form solutions for optimizing neural networks underscores the need to better understand their behavior in practice during training, which is rapidly becoming a computational bottleneck. One approach to address this issue is to employ mathematical abstractions, as seen with neural tangent kernels (NTK) (Jacot et al., 2018; Chizat et al., 2020) and mean-field analysis (Mei et al., 2018; Chen et al., 2024), among others (Abbe et al., 2022; Ahn et al., 2024; Cabannes et al., 2024c). Unfortunately, obtaining formal, rigorous results often requires simplifications that deviate from practical implementations (Bietti et al., 2023; Cabannes et al., 2024c; Boix-Adsera et al., 2023; Tian et al., 2023; Nichani et al., 2024; Edelman et al., 2024; Cabannes et al., 2024b; Ahn et al., 2023; Geshkovski et al., 2023; Jelassi et al., 2022; Mahankali et al., 2024). This type of analysis can overlook crucial details that significantly alter the training dynamics. In our work, we try to remain as faithful as possible to the original Transformer architecture to ensure that our insights bear real-world implications.
2 Controlled Setup
@@
In this section, we describe our controlled setup, focusing on the data and architecture, along with our visualization tools to better understand the training dynamics.
@@ 2.1 Sparse Modular Addition
In terms of data, we concentrate on a straightforward mathematical task. We draw inspiration from the sparse parity problem (Barak et al., 2023; Abbe et al., 2023) and the modular addition problem (Power et al., 2022; Nanda et al., 2023) to formulate the sparse modular addition problem. This problem is characterized by the following parameters:
@@
$\begin{array}[]{cccc}f^{*}:&(\mathbb{F}_{p})^{L}&\to&\mathbb{F}_{p}\\ &x=(x_{1},\ldots,x_{L})&\mapsto&y=\sum_{t\in[k]}x_{t}.\end{array}$
@@
2.2 Transformer Architecture
Our study is made with a transformer model that adapts the original implementation by Vaswani et al. (2023) to our task.
@@ 2.2 Transformer Architecture
The sparse modular addition problem is inherently discrete. To handle it with a differentiable architecture, we need to embed it in a continuous space ${\mathbb{R}}^{d}$. We first embed each token with both semantic and positional information. Given a learnable token embedding $E:{\mathbb{F}}_{p}\to{\mathbb{R}}^{d}$, and a learnable position embedding $P:[L]\to{\mathbb{R}}^{d}$, a sentence in token space is lifted to a sentence in embedding space through the following embedding operation that applies on each $x_{t}$ for $t\in[L]$
$z_{t}:=Z(x_{t},t):=\frac{E(x_{t})+P(t)}{\left\|E(x_{t})+P(t)\right\|}.$
This type of embedding, known as absolute position embedding, incorporates both the token’s semantic meaning and its position within the sequence. The normalization, which projects each element of the sentence onto the sphere, is known as RMS-norm (Zhang & Sennrich, 2019).
Next, a cross-attention mechanism is applied to the sequence of normalized embeddings $(z_{t})$ to aggregate them into a single sentence embedding $\xi\in{\mathbb{R}}^{d}$. It utilizes a query vector $q\in{\mathbb{R}}^{d}$, and a value matrix $V\in{\mathbb{R}}^{d\times d}$. Denoting $z=(z_{t})\in{\mathbb{R}}^{d\times L}$, the sentence embedding can be expressed in matrix form as
@@
where the softmax operation is defined as a mapping from ${\mathbb{R}}^{L}$ to the simplex $\Delta_{L}$ in ${\mathbb{R}}^{L}$. Specifically, the $s$-th component of $\operatorname{softmax}((a_{t}))$ is proportional to $\exp(a_{s})$, formally expressed as $\operatorname{softmax}((a_{t})_{t\in[L]})=(\exp(a_{s})/\sum_{t\in[L]}\exp(a_{t}))_{s\in[L]}$. Compared to the attention mechanism in Vaswani et al. (2023), we omit both the key and output matrices, which would act as extra parameters that do not increase the expressivity of our model since $z$ can be set to anything thanks to $E$ and $P$.
@@
$\psi:=\xi+\sum_{i\in[h]}u_{i}\cdot\sigma\left(\frac{w_{i}^{\top}\xi}{\|\xi\|}\right)\in\mathbb{R}^{d},$
where $\sigma:\mathbb{R}\to\mathbb{R}$ is a non-linear function, chosen to be $\sigma(x)=x\varphi(x)$ with $\varphi$ being the Gaussian cumulative function.
The function $\sigma$ is known as the GELU activation. The final embedding vector $\zeta$ is decoded back to token space based on how it aligns with the respective token embeddings which amounts to using weight-tying in the implementation (Press & Wolf, 2017). More specifically, the model’s logits write
$\zeta=(E(v)^{\top}\psi)_{v\in\mathbb{F}_{p}}\in\mathbb{R}^{p}.$
A softmax layer converts these alignments into a probability vector $\mu=\operatorname{softmax}(\zeta)\in\Delta_{p}$ over the different output classes. Abstracting all the learnable weights into a single parameter $\theta$, our architecture takes as input a sentence $x\in(\mathbb{F}_{p})^{L}$ and outputs the probability $\hat{p}_{\theta}(y|x)=\mu_{y}$. The parameters of our model are optimized by minimizing the cross-entropy loss defined as:
$\mathcal{L}(\theta):=\mathbb{E}_{(x,y)\sim p}\left[-\log(\hat{p}_{\theta}(y|x))\right],$
where $p$ is a training distribution, typically chosen as the counting measure over some training data. This loss is a proxy for the measure we aim to optimize, i.e., the accuracy [^1] In practice, the cross-entropy is a differentiable proxy of the error, i.e., the $0-1$ loss which is the counterpart of the accuracy., defined as
$\mathcal{A}(\theta):=\mathbb{E}_{(x,y)\sim q}\left[\mathbbm{1}_{\{y=\operatorname*{arg\,max}_{v}\hat{p}_{\theta}(v|x)\}}\right],$
with $q$ denoting a data distribution, typically chosen as the counting measure over some testing data. The implementation details are given in Appendix A.
2.3 Visual Sandbox
To visualize everything occurring within our model along the training, we embed all computations in the plan by taking $d=2$. Note that we also validate our insights in larger models with embeddings size $d>2$ in Appendix B. Our codebase is designed to generate videos of the training dynamics, tracking several key aspects. We notably track the following.
@@ 2.3 Visual Sandbox
Figure 1: Clustering Head. Implementation of an idealized circuit that captures the invariants of the problem.
@@
Transform level lines. We visualize the mapping from sentence embedding $\xi$ to their associated learned probabilities $\hat{p}_{\theta}(y|x)$. We also plot the sentence embedding on the same plot to better understand the level line changes.